OpenStudy (anonymous):

How do I determine the slope of the line through the points (1,-8) and (-5,-4)? Please help

OpenStudy (anonymous):

y2-y1 / x2-x1

OpenStudy (anonymous):

What is that called? @telltoamit

OpenStudy (anonymous):

its formula for slope

OpenStudy (anonymous):

@telltoamit So 4/-6 ? Is that correct ?

OpenStudy (anonymous):

Thats my slope?

OpenStudy (anonymous):

yepp. jus plug in the values

Directrix (directrix):

(1,-8) and (-5,-4) Think of slope of a line as rise over run or as change in y over change in x. slope= [ (-4) - (-8)] / [(-5) - (1)] = (4)/(-6) = - 2/3

OpenStudy (anonymous):

Thank you!!! @Directrix @telltoamit @Directrix -2/3 thats because I have to simplify right?

Directrix (directrix):

After you learn the slope concept, you may want to check out the online slope calculator at the following link: http://www.basic-mathematics.com/slope-calculator.html

Directrix (directrix):

Yes. -4/6 reduces to -2/3 because -2 and 6 both have a common factor of 2. @desibaybee

OpenStudy (anonymous):

@Directrix THANK YOU SO MUCH :D I forgot you have to simplify lol Almost left -4/6 as my answer whoops!